# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.4 Proof of Theorem 3
$\left|\frac{1}{6}\sum_{l=1}^{T}\sum_{k=1}^{T}\sum_{v=1}^{T}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\left(\boldsymbol{z}_{d,l}-\boldsymbol{z}_{d,0,l}\right)\left(\boldsymbol{z}_{d,k}-\boldsymbol{z}_{d,0,k}\right)\left(\boldsymbol{z}_{d,v}-\boldsymbol{z}_{d,0,v}\right)\right|$
$\leq C\beta^{2}T^{-3/2}\log^{3}(T),$
$\left|\frac{1}{6}\sum_{l=1}^{T}\sum_{k=1}^{T}\sum_{v=1}^{T}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d+1,0}+\delta\left(\boldsymbol{z}_{d+1}-\boldsymbol{z}_{d,0}\right)\right)\left(\boldsymbol{z}_{d+1,l}-\boldsymbol{z}_{d,0,l}\right)\left(\boldsymbol{z}_{d+1,k}-\boldsymbol{z}_{d,0,k}\right)\left(\boldsymbol{z}_{d+1,v}-\boldsymbol{z}_{d,0,v}\right)\right|$
$\leq C\beta^{2}T^{-3/2}\log^{3}(T),$
@@
$\left|\mathbb{E}\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)\right\}-\mathbb{E}\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\right\}\right|\leq C\beta^{2}T^{-3/2}\log^{3}(T).$
@@
$\left|\mathbb{E}\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)^{2}\right\}-\mathbb{E}\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)^{2}\right\}\right|\leq C\beta^{2}T^{-3/2}\log^{3}(T),$
@@
$\left|\mathbb{E}\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\left(V_{d}-V_{d,0}\right)\right\}-\mathbb{E}\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\right\}\right|$
$\quad\leq C\beta^{2}T^{-3/2}\log^{3}(T).$
Hence, we have $\mathbb{E}\left(\left|V_{d}-V_{d,0}\right|^{3}\right)=O\left(T^{-3/2}\log^{3}(T)\right)$. For $\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{3}\right)$, we first calculate $\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$, then it’s easy to get the order for 3-order term.
$\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$ $=\mathbb{E}\left(\frac{N\omega_{T}^{-1}\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+N\omega_{T}^{-1}\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}}{\sigma_{SS}}\right)^{4}$
$=\frac{N^{4}\omega_{T}^{-4}}{4\left(\operatorname{tr}\left(\mathbf{R}^{2}\right)\right)^{2}}\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}.$
@@
$(i)$ $=\mathbb{E}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4},(ii)=\mathbb{E}\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{3},$
$(iii)$ $=\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}\boldsymbol{Y}_{i}^{\top}h_{i}h_{d}\boldsymbol{Y}_{d}\right)^{2},(iv)=\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{3}\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d},$
$(v)$ $=\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}.$ (12)
Since $\mathbb{E}\boldsymbol{Y}_{i}=\mathbb{E}\boldsymbol{\xi}_{i}=0$, we easily find that (A.4)-(ii)(iv) equal to 0. Next we can get the following equations for (A.4)-(iii) after some straightforward calculations.
$\quad\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}$
$=\mathbb{E}\left[\mathbb{E}\left[\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\mid\boldsymbol{Y}_{d}\right]\right]$
$=\mathbb{E}\left[\sum_{i=1}^{d-1}\sum_{j=d+1}^{T}h_{i}^{2}h_{j}^{2}h_{d}^{4}\left(\boldsymbol{Y}_{d}^{\top}\boldsymbol{\Sigma}_{u}\boldsymbol{Y}_{d}\right)^{2}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{d-1}\sum_{j=d+1}^{T}h_{i}^{2}h_{j}^{2}h_{d}^{4}\left(\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)^{\top}\boldsymbol{\Sigma}_{u}^{2}\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)\right)^{2}\right]$
@@
$\mathbb{E}\left(\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{X}\right)^{2}$ $=2\operatorname{tr}\left(\mathbf{A}^{2}\right)+\operatorname{tr}^{2}(\mathbf{A})\leq 2\operatorname{tr}^{2}\left(\mathbf{A}\right),$
$\mathbb{E}\left(\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{Y}\right)^{4}$ $=\mathbb{E}\left[\mathbb{E}\left[\left(\boldsymbol{Y}^{\top}\mathbf{A}\boldsymbol{X}\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{Y}\right)^{2}\mid\boldsymbol{X}\right]\right]\leq 2\mathbb{E}\left[\operatorname{tr}^{2}\left(\mathbf{A}\boldsymbol{X}\boldsymbol{X}^{\top}\mathbf{A}\right)\right]$
$=2\mathbb{E}\left[\left(\boldsymbol{X}^{\top}\mathbf{A}^{2}\boldsymbol{X}\right)^{2}\right]\leq 4\operatorname{tr}^{2}\left(\mathbf{A}^{2}\right).$
@@
$\mathbb{E}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}$ $=\mathbb{E}\left(\left(\frac{1}{\sqrt{\sum_{i=d+1}^{T}h_{i}^{2}}}\boldsymbol{\Sigma}_{u}^{-1/2}\sum_{i=d+1}^{T}h_{i}\boldsymbol{Y}_{i}\right)^{\top}\left(\sqrt{\sum_{i=d+1}^{T}h_{i}^{2}h_{d}^{2}}\boldsymbol{\Sigma}_{u}\right)\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)\right)^{4}$
@@
$\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}\leq\left(\sum_{i=1}^{d-1}h_{i}^{2}\right)^{2}h_{d}^{4}O\{\operatorname{tr}^{2}\left(\boldsymbol{\Sigma}_{u}^{2}\right)\}$
@@
$\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$
$=$ $\frac{N^{4}\omega_{T}^{-4}}{4\left\{\operatorname{tr}\left({\bf R}^{2}\right)\right\}^{2}}\mathbb{E}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}$
@@
$\sum_{d=1}^{T}\mathbb{E}\left|W_{d}-W_{d,0}\right|^{3}\leq\sum_{d=1}^{T}\left(\mathbb{E}\left(W_{d}-W_{d,0}\right)^{4}\right)^{3/4}\leq C^{\prime}T^{-1/2},$
@@
$\sum_{d=1}^{T}\left|\mathbb{E}\left\{f\left(W_{d},V_{d}\right)\right\}-\mathbb{E}\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\right|\leq C\beta^{2}T^{-1/2}\log^{3}(T)+C^{\prime}T^{-1/2}\rightarrow 0,$
@@ A.5 Proof of Theorem 4
$T^{1/2}\mathbf{D}^{-1/2}\hat{\boldsymbol{\theta}}=T^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{T}(\boldsymbol{U}_{i}+(1-\eta){\boldsymbol{\alpha}})+C_{T}.$
@@
$\left\|N^{1/2}\sum_{t=1}^{T}\omega_{T}^{-1/2}h_{t}\boldsymbol{Y}_{t}\right\|^{2}\text{ and }\left\|\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=1}^{T}\left(\boldsymbol{Y}_{i}+\zeta_{1}\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right)\right\|_{\infty}^{2},$
@@
are asymptotic independent, where $\left\{\boldsymbol{Y}_{1},\boldsymbol{Y}_{2},\cdots,\boldsymbol{Y}_{T}\right\}$ are sample from $N\left(0,\mathbb{E}\boldsymbol{U}_{1}^{\top}\boldsymbol{U}_{1}\right)$ and $\boldsymbol{Y}_{t}=(y_{1t},\dots,y_{Nt})$. Denote $N^{1/2}\sum_{t=1}^{T}\omega_{T}^{-1/2}h_{t}\boldsymbol{Y}_{t}:=\boldsymbol{\varphi}=\left(\varphi_{1},\cdots,\varphi_{N}\right)^{\top},\boldsymbol{\varphi}_{\mathcal{A}}=\left(\varphi_{j_{1}},\cdots,\varphi_{j_{d}}\right)^{\top}$, and $\boldsymbol{\varphi}_{\mathcal{A}^{c}}=\left(\varphi_{j_{d+1}},\cdots,\varphi_{j_{N}}\right)^{\top}$, where $\mathcal{A}=\left\{j_{1},j_{2},\cdots,j_{d}\right\}$ is the set of $\{i:\alpha_{i}\neq 0\}$ and $d=|\mathcal{A}|$. Then, $S=\|\boldsymbol{\varphi}\|^{2}=\left\|\boldsymbol{\varphi}_{\mathcal{A}}\right\|^{2}+\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$,
@@
$M=$ $\left\|\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{t=1}^{T}\boldsymbol{Y}_{t}+\sqrt{E(r_{t}^{2})T^{-1}}\zeta_{1}\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right\|_{\infty}$
@@
$=$ $\max\left\{\max_{i\in\mathcal{A}}\left(\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{t=1}^{T}{y}_{it}+\sqrt{E(r_{t}^{2})T^{-1}}\zeta_{1}\mathbf{D}^{-1/2}(1-\eta)\boldsymbol{\alpha}\right),\max_{i\in\mathcal{A}^{c}}\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{t=1}^{T}{y}_{it}\right\}.$
@@
From the proof of Theorem 6, we know that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ and $\max_{i\in\mathcal{A}^{c}}\sqrt{E(r_{t}^{2})/T}\sum_{t=1}^{T}{y}_{it}$ are asymptotically independent. Hence, it suffice to show that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ is asymptotically independent with $\max_{i\in\mathcal{A}}\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{t=1}^{T}{y}_{it}$. Because $E(r_{t}^{2})$ is non-random, hence, we just need to prove that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ is asymptotically independent with $\max_{i\in\mathcal{A}}\sqrt{N/T}\sum_{t=1}^{T}{y}_{it}$.

### v2

@@ A.4 Proof of Theorem 3
$\left|\frac{1}{6}\sum_{l=1}^{N}\sum_{k=1}^{N}\sum_{v=1}^{N}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\left(\boldsymbol{z}_{d,l}-\boldsymbol{z}_{d,0,l}\right)\left(\boldsymbol{z}_{d,k}-\boldsymbol{z}_{d,0,k}\right)\left(\boldsymbol{z}_{d,v}-\boldsymbol{z}_{d,0,v}\right)\right|$
$\leq C\beta^{2}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N),$
$\Big{|}\frac{1}{6}\sum_{l=1}^{N}\sum_{k=1}^{N}\sum_{v=1}^{N}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d+1,0}+\delta\left(\boldsymbol{z}_{d+1}-\boldsymbol{z}_{d,0}\right)\right)\left(\boldsymbol{z}_{d+1,l}-\boldsymbol{z}_{d,0,l}\right)\left(\boldsymbol{z}_{d+1,k}-\boldsymbol{z}_{d,0,k}\right)$
×$\quad\quad\times\left(\boldsymbol{z}_{d+1,v}-\boldsymbol{z}_{d,0,v}\right)\Big{|}$
$\leq C\beta^{2}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N),$
@@
$\left|E\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)\mid\mathcal{F}_{fr}\right\}-E\left\{f_{2}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\mid\mathcal{F}_{fr}\right\}\right|$
$\leq$ $C\beta^{2}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N).$
@@
$\left|E\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d}-V_{d,0}\right)^{2}\mid\mathcal{F}_{fr}\right\}-E\left\{f_{22}\left(W_{d,0},V_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)^{2}\mid\mathcal{F}_{fr}\right\}\right|$
$\leq$ $C\beta^{2}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N),$
@@
$|E\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\left(V_{d}-V_{d,0}\right)\mid\mathcal{F}_{fr}\right\}$
$-E\left\{f_{12}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\left(V_{d+1}-V_{d,0}\right)\mid\mathcal{F}_{fr}\right\}|$
$\quad\leq C\beta^{2}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N).$
Hence, we have $E\left(\left|V_{d}-V_{d,0}\right|^{3}\mid\mathcal{F}_{fr}\right)=O\Big{\{}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N)\Big{\}}$.
For $E\left(\left|W_{d}-W_{d,0}\right|^{3}\mid\mathcal{F}_{fr}\right)$, we first calculate $E\left(\left(W_{d}-W_{d,0}\right)^{4}\mid\mathcal{F}_{fr}\right)$, then it’s easy to get the order for 3-order term.
$E\left(\left(W_{d}-W_{d,0}\right)^{4}\mid\mathcal{F}_{fr}\right)$ $=E\left(\frac{N\omega_{T}^{-1}\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+N\omega_{T}^{-1}\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}}{\sigma_{SS}}\right)^{4}$
$=\frac{N^{4}\omega_{T}^{-4}}{4\left(\operatorname{tr}\left(\mathbf{R}^{2}\right)\right)^{2}}E\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}.$
@@
$(i)$ $=E_{1}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4},(ii)=E_{1}\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{3},$
$(iii)$ $=E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}\boldsymbol{Y}_{i}^{\top}h_{i}h_{d}\boldsymbol{Y}_{d}\right)^{2},$
$(iv)$ $=E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{3}\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d},(v)=E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4},$ (10)
where $E_{1}$ means conditional expectation given $\mathcal{F}_{fr}$. Since $E\boldsymbol{Y}_{i}=E\boldsymbol{\xi}_{i}=0$, we easily find that (A.4)-(ii)(iv) equal to 0. Next we can get the following equations for (A.4)-(iii) after some straightforward calculations.
$\quad E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}$
$=E_{1}\left[E_{1}\left[\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{2}\mid\boldsymbol{Y}_{d}\right]\right]$
$=E_{1}\left[\sum_{i=1}^{d-1}\sum_{j=d+1}^{T}h_{i}^{2}h_{j}^{2}h_{d}^{4}\left(\boldsymbol{Y}_{d}^{\top}\boldsymbol{\Sigma}_{u}\boldsymbol{Y}_{d}\right)^{2}\right]$
$=E_{1}\left[\sum_{i=1}^{d-1}\sum_{j=d+1}^{T}h_{i}^{2}h_{j}^{2}h_{d}^{4}\left(\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)^{\top}\boldsymbol{\Sigma}_{u}^{2}\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)\right)^{2}\right]$
@@
$E_{1}\left(\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{X}\right)^{2}$ $=2\operatorname{tr}\left(\mathbf{A}^{2}\right)+\operatorname{tr}^{2}(\mathbf{A})\leq 2\operatorname{tr}^{2}\left(\mathbf{A}\right),$
$E_{1}\left(\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{Y}\right)^{4}$ $=E_{1}\left[E_{1}\left[\left(\boldsymbol{Y}^{\top}\mathbf{A}\boldsymbol{X}\boldsymbol{X}^{\top}\mathbf{A}\boldsymbol{Y}\right)^{2}\mid\boldsymbol{X}\right]\right]\leq 2E_{1}\left[\operatorname{tr}^{2}\left(\mathbf{A}\boldsymbol{X}\boldsymbol{X}^{\top}\mathbf{A}\right)\right]$
$=2E_{1}\left[\left(\boldsymbol{X}^{\top}\mathbf{A}^{2}\boldsymbol{X}\right)^{2}\right]\leq 4\operatorname{tr}^{2}\left(\mathbf{A}^{2}\right).$
@@
$E_{1}\left(\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}$
$=$ $E_{1}\left(\left(\frac{1}{\sqrt{\sum_{i=d+1}^{T}h_{i}^{2}}}\boldsymbol{\Sigma}_{u}^{-1/2}\sum_{i=d+1}^{T}h_{i}\boldsymbol{Y}_{i}\right)^{\top}\left(\sqrt{\sum_{i=d+1}^{T}h_{i}^{2}h_{d}^{2}}\boldsymbol{\Sigma}_{u}\right)\left(\boldsymbol{\Sigma}_{u}^{-1/2}\boldsymbol{Y}_{d}\right)\right)^{4}$
@@
$E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}\leq\left(\sum_{i=1}^{d-1}h_{i}^{2}\right)^{2}h_{d}^{4}O\{\operatorname{tr}^{2}\left(\boldsymbol{\Sigma}_{u}^{2}\right)\}$
@@
$E_{1}\left(\left(W_{d}-W_{d,0}\right)^{4}\right)$
$=$ $\frac{N^{4}\omega_{T}^{-4}}{4\left\{\operatorname{tr}\left({\bf R}^{2}\right)\right\}^{2}}E_{1}\left(\sum_{i=1}^{d-1}h_{i}h_{d}\boldsymbol{\xi}_{i}^{\top}\boldsymbol{Y}_{d}+\sum_{i=d+1}^{T}h_{i}h_{d}\boldsymbol{Y}_{i}^{\top}\boldsymbol{Y}_{d}\right)^{4}$
@@
$\sum_{d=1}^{T}E_{1}\left|W_{d}-W_{d,0}\right|^{3}\leq\sum_{d=1}^{T}\left(E_{1}\left(W_{d}-W_{d,0}\right)^{4}\right)^{3/4}\leq C^{\prime}T^{-1/2},$
@@
$\sum_{d=1}^{T}E\mid E\left\{f\left(W_{d},V_{d}\right)|\mathcal{F}_{fr}\right\}-E\left\{f\left(W_{d+1},V_{d+1}|\mathcal{F}_{fr}\right)\right\}\mid$
$\leq$ $C\beta^{2}E\sum_{d=1}^{T}|1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd}|^{3}T^{-3/2}\log^{3}(N)+C^{\prime}T^{-1/2}$
$\leq$ $C\beta^{2}O(1)T^{-1/2}\log^{3}(N)+C^{\prime}T^{-1/2}\rightarrow 0,$
@@ A.5 Proof of Theorem 4
$T^{1/2}\mathbf{D}^{-1/2}\hat{\boldsymbol{\theta}}=T^{-1/2}\zeta_{1}^{-1}\sum_{i=1}^{T}\Big{\{}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})\boldsymbol{U}_{i}+\omega{\boldsymbol{\alpha}}\Big{\}}+C_{T}.$
@@
$\left\|N^{1/2}\sum_{t=1}^{T}\omega_{T}^{-1/2}h_{t}\boldsymbol{Y}_{t}\right\|^{2}\text{ and }\left\|\sqrt{\frac{N}{T}}\sum_{i=1}^{T}\Big{\{}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})\boldsymbol{Y}_{i}+\zeta_{1}\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\Big{\}}\right\|_{\infty}^{2},$
@@
are asymptotic independent, where $\left\{\boldsymbol{Y}_{1},\boldsymbol{Y}_{2},\cdots,\boldsymbol{Y}_{T}\right\}$ are sample from $N\left(0,E\boldsymbol{U}_{1}^{\top}\boldsymbol{U}_{1}\right)$ and $\boldsymbol{Y}_{t}=(y_{1t},\dots,y_{Nt})$. Denote $N^{1/2}\sum_{t=1}^{T}\omega_{T}^{-1/2}h_{t}\boldsymbol{Y}_{t}:=\boldsymbol{\varphi}=\left(\varphi_{1},\cdots,\varphi_{N}\right)^{\top},\boldsymbol{\varphi}_{\mathcal{A}}=\left(\varphi_{j_{1}},\cdots,\varphi_{j_{d}}\right)^{\top}$, and $\boldsymbol{\varphi}_{\mathcal{A}^{c}}=\left(\varphi_{j_{d+1}},\cdots,\varphi_{j_{N}}\right)^{\top}$, where $\mathcal{A}=\left\{j_{1},j_{2},\cdots,j_{d}\right\}$ is the set of $\{i:\alpha_{i}\neq 0\}$ and $d=|\mathcal{A}|$. Then, $S=\|\boldsymbol{\varphi}\|^{2}=\left\|\boldsymbol{\varphi}_{\mathcal{A}}\right\|^{2}+\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$,
@@
$M=$ $\left\|\sqrt{\frac{N}{T}}\sum_{t=1}^{T}\boldsymbol{Y}_{t}+\sqrt{NT^{-1}}\zeta_{1}\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right\|_{\infty}$
@@
$=$ $\max\bigg{\{}\max_{i\in\mathcal{A}}\left(\sqrt{\frac{N}{T}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){y}_{it}+\sqrt{NT^{-1}}\zeta_{1}\mathbf{D}^{-1/2}\omega\boldsymbol{\alpha}\right),$
$\quad\quad\quad\max_{i\in\mathcal{A}^{c}}\sqrt{\frac{N}{T}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){y}_{it}\bigg{\}}.$
@@
From the proof of Theorem 6, we know that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ and
$\max_{i\in\mathcal{A}^{c}}\sqrt{NT^{-1}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){y}_{it}$
are asymptotically independent. Hence, it suffice to show that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ is asymptotically independent with $\max_{i\in\mathcal{A}}\sqrt{NT^{-1}}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){y}_{it}$. Hence, we just need to prove that $\left\|\boldsymbol{\varphi}_{\mathcal{A}^{c}}\right\|^{2}$ is asymptotically independent with $\max_{i\in\mathcal{A}}\sqrt{N/T}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){y}_{it}$.